Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability questions

Expert replies
by [email protected] » Tue Aug 12, 2014 3:38 pm
Hi guys,

Really stuck on these two questions. I paused my mock paper to spend 15 mins working each of them out and still didn't get anywhere! Would appreciate any guidance...

Thanks
Shakira


Image

Image
Join the discussion
Source: — Problem Solving |

by GMATinsight » Tue Aug 12, 2014 6:19 pm
[email protected] wrote:Hi guys,

Really stuck on these two questions. I paused my mock paper to spend 15 mins working each of them out and still didn't get anywhere! Would appreciate any guidance...

Thanks
Shakira


Image
SOLUTION TO QUESTION 2

Favorable Out comes = 3 Pies and 2 cakes = 5C3 x 6C2 = 10 x 15 = 150

Total Outcomes = any 5 out of 11 pieces = 11C5 = 462

Probability = 150/462 = 75/231 = 25/77

Answer: Option B
"GMATinsight"Bhoopendra Singh & Sushma Jha
Most Comprehensive and Affordable Video Course 2000+ CONCEPT Videos and Video Solutions
Whatsapp/Mobile: +91-9999687183 l [email protected]
Contact for One-on-One FREE ONLINE DEMO Class Call/e-mail
Most Efficient and affordable One-On-One Private tutoring fee - US$40-50 per hour
Join the discussion

by GMATGuruNY » Tue Aug 12, 2014 8:15 pm
Each of 4 bags contains 25 blue disks, 25 green disks, 25 orange disks, 25 yellow disks, and nothing else. If one disk is chosen at random from each of the four bags, what is the probability that the number of blue disks chosen will be no less than 1 and no greater than 3?
a. 1/16
b. 41/128
c. 87/128
d. 225/256
e.255/256

OA is C
Here:
A GOOD outcome is selecting 1, 2, or 3 blue disks.
A BAD outcome is selecting no blue disks or 4 blue disks.

P(good outcome) = 1 - P(bad outcome).

In each bag, 1/4 of the disks are blue, while 3/4 are not blue.
P(4 non-blue disks) = 3/4 * 3/4 * 3/4 * 3/4 = 81/256.
P(4 blue disks) = 1/4 * 1/4 * 1/4 * 1/4 = 1/256.
Thus:
P(good outcome) = 1 - (81/256 + 1/256) = 174/256 = 87/128.

The correct answer is C.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by [email protected] » Wed Aug 13, 2014 3:03 pm
GMATinsight wrote:
[email protected] wrote:Hi guys,

Really stuck on these two questions. I paused my mock paper to spend 15 mins working each of them out and still didn't get anywhere! Would appreciate any guidance...

Thanks
Shakira


Image
SOLUTION TO QUESTION 2

Favorable Out comes = 3 Pies and 2 cakes = 5C3 x 6C2 = 10 x 15 = 150

Total Outcomes = any 5 out of 11 pieces = 11C5 = 462

Probability = 150/462 = 75/231 = 25/77

Answer: Option B
Thank you! I'm not sure why I didn't use combinatorics to begin with.
Join the discussion

by [email protected] » Wed Aug 13, 2014 3:07 pm
GMATGuruNY wrote:
Each of 4 bags contains 25 blue disks, 25 green disks, 25 orange disks, 25 yellow disks, and nothing else. If one disk is chosen at random from each of the four bags, what is the probability that the number of blue disks chosen will be no less than 1 and no greater than 3?
a. 1/16
b. 41/128
c. 87/128
d. 225/256
e.255/256

OA is C
Here:
A GOOD outcome is selecting 1, 2, or 3 blue disks.
A BAD outcome is selecting no blue disks or 4 blue disks.

P(good outcome) = 1 - P(bad outcome).

In each bag, 1/4 of the disks are blue, while 3/4 are not blue.
P(4 non-blue disks) = 3/4 * 3/4 * 3/4 * 3/4 = 81/256.
P(4 blue disks) = 1/4 * 1/4 * 1/4 * 1/4 = 1/256.
Thus:
P(good outcome) = 1 - (81/256 + 1/256) = 174/256 = 87/128.

The correct answer is C.
Thank you! Much appreciated
Join the discussion

by shrivats » Wed Aug 13, 2014 7:32 pm
For the first question:

no less than 1 blue disk means, it cannot be zero
no more than 3 means, it canot be all four.

P( zero blue) = (3/4)* (3/4)*(3/4)*(3/4) because 75/100 is 3/4
P(all 4 bue) = (25/100)^4

required probability is = 1 - [ (3/4)^4 * (1/4)^4

= 1 - [ (81/256 + 1/256) ]
= 1 - [ 82 /256 ]
= 174/256 = 87/128
Join the discussion

by shrivats » Wed Aug 13, 2014 7:37 pm
The second is straight forward.
there are 5 pies and 6 cakes

total of 5 items selected out of 11, 3 of which are pies, implies are 2 cakes.

so the probability is = 5c3 * 6c2
----------- = 150/462 = 25/77
11c5
Join the discussion